You have been hired as a credit and analyst for the brand new Chemical Bank. Below is a set of financial statements for a local company who is requesting a loan for $750,000.
Prepare and submitted a typed, written report detailing your findings which support whether you will or will not grant the loan which is requested. All calculations and work need to be included in the report. The analysis should be about 2 pages in length.
ACC 122 Final Case Study-Financial Statement Analysis Smith Company Comparative Income Statements 2019 2018 2017 2016 2015 2014 2013 1950 1700 1500 140012001110 920 1246 1032 752 Gross Profit Operating Exp Interest Exp 20 25 24 270 304 230 118 216 146 374 412 428 100 Net Income 257 372 246 176 Comparative Balance Sheets A/R net Inventory Other Current Assets 000 850 600 450318 302 216 1838 1364 12041032936 810 615 14 Long Term Invest PPE, net Total Assets 16881264 0392954 5973 5278 4617 3482 3080 2494 1868 2667 1832 Liab and Equity Current Liab 1300 1030 718630 600 450 250 LT Liab 1459 1291 1251698 698 728 500 Common Stock APIC Ret Earnings Total Equity 1000 1000 1000 850 850 650650 250 250 250 170 170 15o 150 1964 8214 2957 26482154 1782 1316 1118 1707 1398 1134 516 Total Liab and Equity 5973 5278 4617 3482 3080 2494 1868
Answer #1

Before granting loan of Rs $750,000 to a local company there are certain financial ratios to be determined to assess whether the local company is eligible for loan amount of $750,000. After considering their financial statements and relevant financial ratios, we can take decision whether to grant them loan of $750,000. As a Credit analyst of the brand new Chemical Bank it is important to check current cash flows as well as future cash flow from the business so as to check creditworthiness of the company.

  • Net Present Value of the future Cash Flows (NPV Analysis)- Taking year 2013 as the base year. We need to calculate the Net Present Value (NPV) of the future cash flows coming in the business. Because present value of the future cash flows will be less due to time value of money principle. To discount the future cash flows we can take discounting factor as 10% (Assumption)

NPV of cash flows = 920/(1.10)1+ 1110/(1.10)2 + 1200/(1.10)3 +1400/(1.10)4 + 1500/(1.10)5 +1700/(1.10)6 +1950/(1.10)7

Financial Ratios to ascertain before grant of $750,000 loan

  • Debt- Equity Ratio: It is important to assess that out of all sources of funding or capital raised by the company. How much percentage of the capital is debt-financed. Company having high debt-equity ratio is considered more risky to lend as compared to company with low debt-equity ratio. Debt-Equity Ratio for Year 2019 i.e. 2759/3214 = 0.85
  • Current Ratio - This ratio is used to ascertain whether company is able to meet their day to day expenses or in accounting term working capital requirements. Current ratio is calculated by Current Assets/ Current Liabilities. Current Ratio for year 2019 = 2.25 ( which indicates that Current Assets of the company are 2.25 times than the current liabilities for the year 2019
  • Gross Margin Ratio- This ratio indicates gross profit as the percentage of the Sales Revenue. For 2019 year Gross Margin Ratio = 704/ 1950*100 i.e. 36.10%
  • Net Profit Margin- This financial ratio will determine the overall profitability in terms of net profits as percentage of sales revenue during the time period. For year 2019 Net Profit margin ratio = Net Income/ Sales Revenue*100 i.e. 257/1950*100= 13.18% for year 2019.
  • Return on Investment (ROI)- This ratio ascertain the returns on the investments made by the organization. This ratio is one of the key ratio for assessing the financial soundness and business performance of the company.
  • Debt to Assets ratio - This financial leverage ratio determine what amount of total assets are financed by the borrowed capital. Debt to Assets ratio for this company for 2019 year i.e. 2759/5973*100 = 46.2%
  • Debt Service Coverage Ratio- This ratio help banks and other lending financial institution to ascertain company's capacity to pay off loan payments by the net income generated. DSCR can be calculated by dividing Net Income/ Annual debt service (Loan payments) i.e. 257/2759 = 0.093
  • Loan to Value Ratio- Most of the financial institution including bank require collateral in order to grant loan this is known as secured loans. Collateral kept in the bank for lending purpose is generally higher than the requested loan amount as in case of default, lending institution can raise loan amount by selling the collateral. In this case company is asking for unsecured loan as no information is given regarding any mortgage or collateral.

After analyzing the financial statements and financial ratios of the company, we conclude that company is performing consistently and have financial soundness. However grant of loan will also majorly depend upon the time period horizon, repayment cycle and Interest rate charged by the bank for lending $750,000. Bank can check Interest Coverage ratio before taking decision to grant loan.

As per the given financial statements, Bank should grant loan amount of $750,000 to the company.